OpenStudy (anonymous):

Divide and simplify if possible Show steps √(250x^16)/√(2x)

OpenStudy (anonymous):

please help

OpenStudy (gorv):

((250x^16)/(2x))^(1/2) =((250/2)*(x^16/x))^(1/2) =(125*x^15)^(1/2)

OpenStudy (anonymous):

are those all of the steps?

OpenStudy (gorv):

yeah.....u got it ????

OpenStudy (anonymous):

I think so

OpenStudy (gorv):

(125)^(1/2)*(x^(15))^1/2 =5^3/2 * x^15/2

OpenStudy (anonymous):

more steps?

OpenStudy (gorv):

noooo buddy

OpenStudy (anonymous):

gudjdj